# What makes math problems hard for reinforcement learning: a case study

### v1

@@ Abstract.
Using a long-standing conjecture from combinatorial group theory, we explore, from multiple angles, the challenges of finding rare instances carrying disproportionately high rewards. Based on lessons learned in the mathematical context defined by the Andrews–Curtis conjecture, we propose algorithmic improvements that can be relevant in other domains with ultra-sparse reward problems. Although our case study can be formulated as a game, its shortest winning sequences are potentially $10^{6}$ or $10^{9}$ times longer than those encountered in chess. In the process of our study, we demonstrate that one of the potential counterexamples due to Akbulut and Kirby, whose status escaped direct mathematical methods for 39 years, is stably AC-trivial.
@@
reinforcement learning, large language models, topological data analysis, automated reasoning, search algorithms
@@ 1. Introduction
In this paper, we begin a systematic implementation of this approach by carefully analyzing the distribution of hardness in potential counterexamples to a long-standing conjecture, the Andrews–Curtis conjecture. As with many other challenging problems, a natural measure of hardness is the number of steps an RL agent needs to take. What makes the Andrews–Curtis conjecture particularly suitable for our study is that it includes numerous examples requiring a hyper-exponential number of steps, providing an effective testbed for exploring the aforementioned questions through the lens of RL, search algorithms, and topological data analysis.
@@
While our primary focus is on exploring the distribution of hardness with an eye toward algorithm development, we also resolve a particularly intriguing open case that has eluded direct mathematical approaches for decades:
Theorem 1.
The following potential counterexample introduced by Akbulut and Kirby [Akbulut--Kirby] is stably AC-trivial:
$\text{AK}(3)=\langle x,y\mid x^{3}=y^{4},xyx=yxy\rangle.$
The proof of this theorem is presented in Subsection 3.5. The rest of the paper is organized as follows. In Section 2, we provide an overview of the Andrews–Curtis conjecture, describing the specific version studied in this work. We then apply classical search algorithms to examples in the Miller–Schupp series in Section 3, where we devise a greedy search algorithm that significantly outperforms the widely used breadth-first search algorithm. In Section 4, we employ reinforcement learning, specifically implementing the Proximal Policy Optimization (PPO) algorithm [schulman2017proximal], and find that while it performs better than breadth-first search, it does not surpass the greedy search algorithm (see Figure 1). Building on these insights, Section 5 presents several ideas for algorithm development, proposing strategies to mitigate the overwhelming complexity faced by an RL agent in challenging problems. In Section 6, we employ topological methods to assess the complexity of presentations. Specifically, we utilize persistent homology to define the isolation value of a presentation. This value is infinite for counterexamples to the AC conjecture and serves as a measure of a presentation’s resistance to being trivialized. Finally, in Section 7, we examine the linguistic structure of balanced presentations using a decoder-only Transformer model, observing distinct clusters within the embedding space corresponding to presentations solved and unsolved by the greedy search algorithm.
@@
(a) Distribution versus $n$
(b) Distribution versus length
Figure 1. A comparison of three algorithms —breadth-first search, greedy search, and Proximal Policy Optimization (PPO)— that we used to search through the space of balanced presentations. The number of presentations of the Miller–Schupp series, $\text{MS}(n,w)$, solved by an algorithm is given on the vertical axis. We compare the performance as a function of $n$ (above) and the length of the presentation (below). Greedy Search consistently outperforms Breadth-First Search and Proximal Policy Optimization.
@@ Acknowledgment
We would like to thank Danil Akhtiamov, Anna Beliakova, Jessica Craven, Michael Douglas, Konstantin Korovin, Alexei Lisitsa, Maksymilian Manko, Ciprian Manolescu, Fabian Ruehle, Josef Urban, and Tony Yue Yu for insightful discussions and comments. We especially want to thank Anna Beliakova for igniting our interest in the Andrews–Curtis conjecture as a framework for exploring problems with long and rare sequences of moves that an RL agent must discover.
@@ 2. Andrews–Curtis conjecture
$\text{AK}(2)$ has length 11 and has been established as AC-trivial [genetic] whereas $\text{AK}(3)$ is the aforementioned presentation with length 13.
@@
In over two decades since the first utilization of search algorithms [genetic, bfs-ac], only unsuccessful attempts have been made to trivialize $\text{AK}(3)$ with different variants of breadth-first search algorithm using an increasing amount of computational resources [Bowman-McCaul, krawiec2016distance, Panteleev-Ushakov]. Notably, [Panteleev-Ushakov] found that no sequence of AC-moves that allows relator lengths to increase up to 20 trivializes $\text{AK}(3)$. This lack of success could be interpreted as suggestive evidence that $\text{AK}(3)$ might be a counterexample to the Andrews–Curtis conjecture. However, recent works by Bridson and Lishak have shown that there exist AC-trivializable balanced presentations, for which the number of AC-moves in a trivializing sequence is bounded below by a superexponential function of the length of the presentation [Bridson, Lishak]. Roughly speaking, for these presentations, if the sum of word lengths is $k$, the number of AC-moves required to trivialize the presentation is at least $\Delta(\lfloor\log_{2}k\rfloor)$ where $\Delta\colon\mathbb{N}\to\mathbb{N}$ is defined recursively as $\Delta(0)=2$ and $\Delta(j)=2^{\Delta(j-1)}$ for $j\geq 1$. In particular, $\Delta(\lfloor\log_{2}(13)\rfloor)=65536$, whereas presentations trivialized by the aforementioned search algorithms have AC sequences of length less than $1000$. While $\text{AK}(3)$ is itself not a member of the family of examples studied by Bridson and Lishak, their findings challenge the view it as a counterexample. Their work also underscores the necessity of employing search methods that are more efficient than breadth-first search.
@@
In this paper, we will consider a variety of computational tools to better understand the properties of balanced presentations of the trivial group. We will test the efficacy of our approaches on a subset of presentations from the Miller–Schupp series of potential counterexamples [Miller--Schupp]:
@@
Here, $n\geq 1$ is a positive integer and $w$ is a word in $x$ and $y$ with zero exponent sum on $x$. For $w_{\star}=y^{-1}x^{-1}yxy$, the presentations $\text{MS}(n,w_{\star})$ are AC-equivalent to the presentations from Akbulut–Kirby series [MMS]. In particular, the presentation
$\text{MS}(n,w)=\langle x,y\mid x^{-1}y^{3}x=y^{4},x=y^{-1}x^{-1}yxy\rangle.$
of length 15 is AC-equivalent to $\text{AK}(3)$.
@@
For two generators, which is the only case we study in this paper, the group generated by these AC-transformations is isomorphic to the group generated by the original AC-transformations.333 The difference lies in how the inversion of a relator is handled: we always follow an inversion by a concatenation, while the original AC-moves allow for standalone inversion moves. The original inversion moves may be retrieved from the new generators as follows. For a given presentation $\langle x_{1},x_{2}\mid r_{1},r_{2}\rangle$, the sequence of moves: $r_{2}\to r_{2}r_{1}$, $r_{1}\to r_{1}r_{2}^{-1}$, $r_{2}\to r_{2}r_{1}$, and $r_{2}\to r_{1}r_{2}r_{1}^{-1}$ results in the presentation $\langle x_{1},x_{2}\mid r_{2}^{-1},r_{1}\rangle$, which is the same as $r_{2}\to r_{2}^{-1}$ up to swapping the two relators. We also enhanced the notion of trivial presentation(s) to include all presentations of length 2: $\{\langle x_{1},x_{2}\mid x_{i}^{a},x_{j}^{b}\rangle\mid i,j=1,2;a,b=\pm 1;i\neq j\}$. The reason for this change is due to its effect on performance in greedy search and reinforcement learning algorithms studied in Section 3 and Section 4. Specifically, the length of a presentation provides a useful signal when searching through the space of presentations with these algorithms. An inversion transformation leaves the length invariant providing no signal to the search process and slowing down the performance of the algorithm significantly. For the rest of the paper we will refer to the new transformations (instead of the original AC-transformations) as “AC-transformations” or “AC-moves”.
@@ 3. Classical Search Algorithms
In this section, we compare the effectiveness of breadth-first and greedy search algorithms to AC-trivialize presentations in the Miller–Schupp series. We find that the latter significantly outperforms the former. Additionally, using the greedy search algorithm, we determine that, in the stable case, $\text{AK}(3)$ is AC-trivial.
@@ 3.1. Breadth-first search
We first recall the breadth-first search algorithm. An iterative implementation of this algorithm, adapted to the problem of the Andrews–Curtis conjecture, is provided in Algorithm 1.
We start with an initial state, which is a balanced presentation we wish to AC-trivialize, and place it in a queue. At each iteration, a state is removed from the queue, and its neighbors are added if they have not already been visited. This process continues until the sought-after state, i.e., a trivial balanced presentation, is found or a maximum number of states $N$ is visited. In our experiments, we set $N=10^{6}$.
Algorithm 1 Breadth-First Search Algorithm
@@ 3.2. Greedy search
The greedy search algorithm, described in Algorithm 2, differs only slightly from the breadth-first search algorithm in implementation. We replace the queue with a priority queue, which stores the states in order determined by a tuple of values $(k,l)$, where $k$ is the length of the presentation and $l$ is the path length between the state and the initial state.
@@
Algorithm 2 Greedy Search Algorithm
@@
3.3. Comparison
We find that greedy-search algorithm outperforms the breadth-first search algorithm in the task of AC-trivializing Miller–Schupp presentations Figure 1. Out of the 1190 presentations in the Miller–Schupp series with $n\leq 7$ and $\operatorname{length}(w)\leq 7$, greedy search solved 533 while BFS solved only 278 presentations. Each algorithm was constrained to visit a maximum of 1 million nodes. The percentage of presentations solved by these algorithms decreases monotonically as a function of $n$. Remarkably, the greedy search was able to solve all presentations with $n=1$ or length less than 14. There are, however, six presentations of length 14 that greedy search could not solve. We verified that four of these,

### v2

# —PREPRINT—

@@ Abstract.
Using a long-standing conjecture from combinatorial group theory, we explore, from multiple perspectives, the challenges of finding rare instances carrying disproportionately high rewards. Based on lessons learned in the context defined by the Andrews–Curtis conjecture, we propose algorithmic enhancements and a topological hardness measure with implications for a broad class of search problems. As part of our study, we also address several open mathematical questions. Notably, we demonstrate the length reducibility of all but two presentations in the Akbulut–Kirby series (1981), and resolve various potential counterexamples in the Miller–Schupp series (1991), including three infinite subfamilies.
@@
Reinforcement learning, Andrews–Curtis conjecture, automated reasoning, search algorithms, topological data analysis, unknot diagrams
@@ 1. Introduction
We begin a systematic implementation of this approach by carefully analyzing the distribution of hardness in potential counterexamples to a long-standing conjecture, the Andrews–Curtis conjecture. This setting offers distinct advantages for algorithm development compared to generic mathematical problems. First, the action space in this setting is finite and small, consisting of only a few basic moves. Second, each instance has an intrinsic measure of complexity—its length— relative to which the number of steps required by an RL agent ranges from linear to superexponential [Bridson, Lishak].
This work formulates the Andrews–Curtis trivialization problem as a reinforcement learning environment with long horizons, sparse rewards, and an intrinsic hardness distribution. It systematically explores this hardness distribution across a broad class of examples using classical search, reinforcement learning, language modeling, and topological data analysis. Additionally, it introduces a principled topological measure applicable to similar path-to-base search problems. Finally, it examines how reinforcement learning agents navigate problems of varying difficulty, leveraging insights into problem hardness to propose novel algorithms that dynamically adapt to challenging tasks.
While our primary focus is on exploring the distribution of hardness to inform algorithm development, we also discover new mathematical results about long-standing potential counterexamples of this conjecture.
A notable family of potential counterexamples, denoted $\text{AK}(n)$, due to Akbulut and Kirby,
$\text{AK}(n)=\langle x,y\mid x^{n}=y^{n+1},xyx=yxy\rangle,\quad n\geq 3$
has been open in mathematics for more than four decades [Akbulut--Kirby]. [^3] The original work included the case $n=2$, but this case was solved approximately two decades later [genetic].The case $n=1$ also gives a balanced presentation of the trivial group, which is easy to solve in a few steps. The length of the presentation $\text{AK}(n)$ is $72n+$ 7, and until now, it was unknown whether this length can even be reduced using AC moves. We find that for every fixed $n$, $\text{AK}(n)$ is AC-equivalent to each element of the following 1-parameter family of presentations parameterized by $k\in\mathbb{Z}$,
$P(n,k)=\langle x,y\mid y^{n-k-1}x^{-1}yx=xyx^{-1}y^{n-k}\ ,\ x=y^{-k}x^{-1}yxy\rangle.$
When $n\geq 5$, this family contains presentations of length less than $72n+$ 7. The shortest presentation occurs in the case $k=n-1$, which gives us the following result, proven as Theorem 4 in the main text.
Theorem A.
For every $n\geq 2$, $\text{AK}(n)$ is AC-equivalent to the presentation
$\langle x,y\mid x^{-1}yx=xyx^{-1}y\ ,\ xy^{n-1}x=yxy\rangle,$
of length $n+11$. This gives a reduction in length of $AK(n)$ for all $n\geq 5$.
Another important family of potential counterexamples, due to Miller and Schupp, is
$\text{MS}(n,w)=\langle x,y\mid x^{-1}y^{n}x=y^{n+1},x=w\rangle,$
where $n>0$ and $w$ is a word with exponent sum zero on $x$. It has been open for more than 25 years [Miller--Schupp].
Theorem B.
The following infinite subfamilies of Miller–Schupp presentations are AC-trivial:
  1. (i)
    $\text{MS}(1,w)$ for all $w$,
  2. (ii)
    $\text{MS}(n,w_{\star})$ for all $n>0$ and $w_{\star}=y^{-1}xyx^{-1}$,
  3. (iii)
    $\text{MS}(2,w_{k})$ for $w_{k}=y^{-k}x^{-1}yxy$ where $k\in\mathbb{Z}$.
Furthermore, for each fixed $n>0$, all members of the 1-parameter family, $\text{MS}(n,w_{k})$, parameterized by $k\in\mathbb{Z}$, are AC-equivalent to each other and to $\text{AK}(n)$.
In addition to these results involving infinite families, respectively proven as Theorems 2, 3 and 9, we also found AC-trivializations of hundreds of Miller–Schupp presentations. Of special note are two hard presentations, which our classical search algorithms could not solve at first, but were solved by a reinforcement learning agent (Theorem 10):
$\langle x,y\mid x^{-1}y^{3}xy^{-4},\ x^{-1}y^{-1}xy^{-3}x^{-1}y^{-1}\rangle,$
$\langle x,y\mid x^{-1}y^{4}xy^{-5},\ x^{-1}yxy^{-1}x^{-1}y^{-3}\rangle.$
Note that the first of these presentations is extremely similar to the following rewriting of $\text{MS}(3,w_{-3})$,
$\text{MS}(3,w_{-3})=\langle x,y\mid x^{-1}y^{3}xy^{-4},x^{-1}y^{-1}xy^{-3}xy^{-1}\rangle.$
with all but one letter being different. The latter presentations is AC-equivalent to $\text{AK}(3)$—perhaps the most important potential counterexample to the Andrews–Curtis conjecture. This shows how very similar-looking presentations can have vastly different hardness properties, emphasizing the need for a systematic understanding of hardness in math problems.
A weaker version of the AC conjecture can be obtained by incorporating additional stable moves. This establishes a direct connection between the conjecture and knot theory, a relationship we explore in Section 9. Motivated by this connection and by the claim in [MMS] that the presentation
(1) $\langle x,y\mid x^{-1}y^{-1}xy^{-1}x^{-1}yxy^{-2}xyx^{-1}y,\ y^{-1}x^{-1}y^{2}x^{-1}y^{-1}xyxy^{-2}x\rangle$
is stably AC-trivial, we apply reinforcement learning to demonstrate its AC equivalence to $\text{AK}(3)$. However, after closer inspection, we discovered a subtle misprint in [MMS, p.10] that had gone unnoticed for more than 20 years. This misprint undermines the claim that (1) is stably AC-trivial.
@@
Outline
In Section 2, we provide an overview of the Andrews–Curtis conjecture.
We then apply classical search algorithms to examples in the Miller–Schupp series in Section 3, where we devise a greedy search algorithm that significantly outperforms the widely used breadth-first search algorithm.
In Section 4, we employ reinforcement learning, specifically the Proximal Policy Optimization (PPO) algorithm [schulman2017proximal], and find that while it performs better than breadth-first search and solves some presentations that are hard for greedy search, it does not strictly surpass the greedy search algorithm.
Building on these insights, Section 5 presents several ideas for algorithm development, proposing strategies to mitigate the overwhelming complexity faced by an RL agent in challenging problems.
In Section 6, we examine the linguistic structure of balanced presentations using a decoder-only Transformer model, observing distinct clusters within the embedding space corresponding to presentations solved and unsolved by the greedy search algorithm.
In Section 7, we introduce a principled measure of the hardness of general path-to-base search problems. We approximate this invariant for the AC and AC trivialization problems showing the consistency of the measure up to an overall scaling.
In Section 8, we analyze the hardness of a presentation by examining how much of it can be predicted from the local structure of the problem. Specifically, we study the subgraph of the AC graph consisting of all presentations connected to a given presentation through a small number of AC moves.
Finally, in Section 9, we study the stable version of the Andrews–Curtis conjecture, and its relationship with knot theory. This relationship has previously been used to give potential counterexamples of the standard Andrews–Curtis conjecture. We prove AC-triviality of a large class of these presentations.
@@ Acknowledgment
We would like to thank Danil Akhtiamov, Anna Beliakova, Jessica Craven, Michael Douglas, Konstantin Korovin, Alexei Lisitsa, Maksymilian Manko, Ciprian Manolescu, Fabian Ruehle, Josef Urban, Richard Wedeen, Tony Yue Yu, and Miriam Lipniacka for insightful discussions and comments. We especially want to thank Anna Beliakova for igniting our interest in the Andrews–Curtis conjecture as a framework for exploring problems with long and rare sequences of moves that an RL agent must discover.
@@ 2. Andrews–Curtis conjecture
$\text{AK}(2)$ has length 11 and has been established to be AC-trivial [genetic], whereas $\text{AK}(3)$ is the aforementioned presentation with length 13.
@@
In over two decades since the first utilization of search algorithms [genetic, bfs-ac], all attempts to trivialize $\text{AK}(3)$ using different variants of breadth-first search algorithms, even with increasing amount of computational resources, have failed [Bowman-McCaul, krawiec2016distance, Panteleev-Ushakov]. Notably, [Panteleev-Ushakov] found that no sequence of AC-moves that allows relator lengths to increase up to 20 trivializes $\text{AK}(3)$. This persistent lack of success could be interpreted as suggestive evidence that $\text{AK}(3)$ might be a counterexample to the Andrews–Curtis conjecture.
However, recent works by Bridson and Lishak have shown that there exist AC-trivializable balanced presentations, for which the number of AC-moves in a trivializing sequence is bounded below by a superexponential function of the length of the presentation [Bridson, Lishak]. Roughly speaking, for these presentations, if the sum of word lengths is $k$, the number of AC-moves required to trivialize the presentation is at least $\Delta(\lfloor\log_{2}k\rfloor)$ where $\Delta\colon\mathbb{N}\to\mathbb{N}$ is defined recursively as $\Delta(0)=2$ and $\Delta(j)=2^{\Delta(j-1)}$ for $j\geq 1$. In particular, $\Delta(\lfloor\log_{2}(13)\rfloor)=65536$, whereas presentations trivialized by the aforementioned search algorithms have AC sequences of length less than $1000$. Although $\text{AK}(3)$ is itself not amongst the presentations studied by Bridson and Lishak, their findings challenge the inclination to view $\text{AK}(3)$ as a counterexample. Their work further highlights the need for novel approaches to study the Andrews-Curtis conjecture.
One approach, which we explore in this paper, is to use compositions of AC-moves instead of AC-moves themselves to solve potential counterexamples of the conjecture. The main motivation here being that while a presentation may require a superexponential number of AC-moves to trivialize, the number of composed moves, which we call “supermoves”, may be much smaller. As an example, consider the following substitution move which comes from [BurnsI],  [^4] A much larger class of “supermoves” called “elementary M-transformations” was studied in [BurnsI]. However, these moves are infinite in number and are hence not suitable for classical search techniques. In Section 5, we propose an alternative method to discover new supermoves using deep learning methods.
Definition 1 (Substitution).
Let $\langle x_{1},\dots,x_{n}\mid r_{1},\dots,r_{i-1},w^{-1}w^{\prime},r_{i+1},\dots r_{n}\rangle$ be a balanced presentation of the trivial group with some words $w$, $w^{\prime}$ in generators. By a sequence of AC-moves, we may replace any occurrence of $w$ in a relator $r_{j}$ (where $j\neq i$) with $w^{\prime}$.
To see how to realize substitutions with AC-moves, consider by means of an example substituting $w^{\prime}$ for $w$ in the relator $w_{1}ww_{2}$: first, conjugate to get $w_{2}w_{1}w$; then multiply by $w^{-1}w^{\prime}$ to write $w_{2}w_{1}w^{\prime}$; and finally, conjugate again to get the required result $w_{1}w^{\prime}w_{2}$. For a relator of the form $w_{1}ww_{2}\dots w_{l-1}ww_{l}$, we may similarly substitute any number of occurences of $w$ with $w^{\prime}$. The number of AC moves packaged into the substitution move is a linear function of the lengths of $w^{\prime}$ and $w_{i}$. We will frequently use this move in Section 3.4 and Section 3.5 to obtain new results pertaining to potential counterexamples of the Andrews-Curtis conjecture.
@@
We also aim to uncover additional insights related to the Andrews-Curtis conjecture. To achieve this, we employ various computational tools in this paper to deepen our understanding of the properties of balanced presentations of the trivial group. We will test the efficacy of our approaches on a subset of presentations from the Miller–Schupp series of potential counterexamples [Miller--Schupp]:
@@
Here, $n\geq 1$ is a positive integer and $w$ is a word in $x$ and $y$ with zero exponent sum on $x$. For $w_{1}=y^{-1}x^{-1}yxy$, the presentations $\text{MS}(n,w_{1})$ are AC-equivalent to the presentations from Akbulut–Kirby series [MMS]. In particular, the presentation
$\text{MS}(3,w_{1})=\langle x,y\mid x^{-1}y^{3}x=y^{4},x=y^{-1}x^{-1}yxy\rangle.$
of length 15 is AC-equivalent to $\text{AK}(3)$.  [^5] We will find, in Section 3.5 of this paper, that $\text{AK}(n)$ is, in fact, AC-equivalent to all members of a 1-parameter family of presentations, $MS(n,w_{k})$, where $w_{k}=y^{-k}x^{-1}yxy$ and $k\in\mathbb{Z}$.
@@
The group generated by these AC-transformations is isomorphic to the group generated by the original AC-transformations.666 The difference lies in how the inversion of a relator is handled: we always follow an inversion by a concatenation, while the original AC-moves allow for standalone inversion moves. The original inversion moves may be retrieved from the new generators through a sequence of transformations. Consider, for example, a two-generator presentation $\langle x_{1},x_{2}\mid r_{1},r_{2}\rangle$. The sequence of moves $r_{2}\to r_{2}r_{1}$, $r_{1}\to r_{1}r_{2}^{-1}$, $r_{2}\to r_{2}r_{1}$, and $r_{2}\to r_{1}r_{2}r_{1}^{-1}$ results in the presentation $\langle x_{1},x_{2}\mid r_{2}^{-1},r_{1}\rangle$, which is the same as $r_{2}\to r_{2}^{-1}$ up to swapping the two relators. We also enhanced the notion of trivial presentation(s) with $n$ generators to include all presentations of length $n$. For $n=2$, these are $\{\langle x_{1},x_{2}\mid x_{i}^{a},x_{j}^{b}\rangle\mid i,j=1,2;a,b=\pm 1;i\neq j\}$. The reason for this change is due to its effect on performance in greedy search and reinforcement learning algorithms studied in Section 3 and Section 4. Specifically, the length of a presentation provides a useful signal when searching through the space of presentations with these algorithms. An inversion transformation leaves the length invariant providing no signal to the search process and slowing down the performance of the algorithm significantly. For the rest of the paper we will refer to the new transformations (instead of the original AC-transformations) as “AC-transformations” or “AC-moves”.
@@ 3. Classical search algorithms
In this section, we compare the effectiveness of breadth-first and greedy search algorithms to AC-trivialize presentations in the Miller–Schupp series. We find that the latter significantly outperforms the former. Based on our empirical observations, we find AC-triviality of infinite subfamilies of the Miller–Schupp series, as well as length reductions of $\text{AK}(n)$.
@@ 3.1. Breadth-first search
In breadth-first search (See Algorithm 1 for pseudocode), we start with an initial state, which is a balanced presentation we aim to AC-trivialize, and place it in a queue. At each iteration, a state is removed from the queue, and its neighbors are added if they have not already been visited. This process continues until the sought-after state, i.e., a trivial balanced presentation, is found or a maximum number of states $N$ is visited. In our experiments, we set $N=10^{6}$.
Algorithm 1 Breadth-First Search
@@ 3.2. Greedy search
Greedy search, described in Algorithm 2, differs only slightly from breadth-first search. We replace the queue with a priority queue, which stores the states in order determined by a tuple of values $(k,l)$, where $k$ is the length of the presentation and $l$ is the path length between the state and the initial state.
@@
Algorithm 2 Greedy Search
@@
3.3. Comparison of performance on Miller–Schupp series
We find that greedy-search outperforms breadth-first search in the task of AC-trivializing Miller–Schupp presentations. Out of the 1190 presentations in the Miller–Schupp series with $n\leq 7$ and $\operatorname{length}(w)\leq 7$, greedy search solved 533 while BFS solved only 278. Each algorithm was constrained to visit a maximum of 1 million nodes. The percentage of presentations solved by these algorithms decreases monotonically as a function of $n$, with greedy search being able to solve all presentations with $n=1$ (see Figure 1). This empirical result encouraged us to prove AC-triviality of the infinite family $\text{MS}(1,w)$, which we give in Section 3.4.
Figure 1. Comparison of greedy and breadth-first search algorithms as a function of $n$. The number of presentations of the Miller–Schupp series, $\text{MS}(n,w)$, solved by an algorithm is given on the vertical axis.
Greedy search was also able to solve all presentations of length less than 14. At length 14, there are six presentations that greedy search could not solve. We verified that four of these,